### v1

# Machine Learning and the Physical Sciences

Abstract
Sparse Autoencoders (SAEs) can efficiently identify candidate monosemantic features from pretrained neural networks for galaxy morphology. We demonstrate this on Euclid Q1 images using both supervised (Zoobot) and new self-supervised (MAE) models. Our publicly released MAE achieves superhuman image reconstruction performance. While a Principal Component Analysis (PCA) on the supervised model primarily identifies features already aligned with the Galaxy Zoo decision tree, SAEs can identify interpretable features outside of this framework. SAE features also show stronger alignment than PCA with Galaxy Zoo labels. Although challenges in interpretability remain, SAEs provide a powerful engine for discovering astrophysical phenomena beyond the confines of human-defined classification.
**footnotetext: The authors contributed equally to this work.
1 Introduction
Galaxy morphology concepts are traditionally encoded in manually-designed taxonomies, e.g., DeVaucouleurs1963 . These may miss concepts that are too rare to be manually detected. New, rare concepts are statistically guaranteed to be represented in the billions of resolved galaxies soon to be imaged by space telescopes like Euclid and Roman (scaramellaEuclidPreparationEuclid2021, ; spergel2015widefieldinfrarredsurveytelescopeastrophysics, ), so we will need a data-driven method to identify these rare concepts. Relatedly, the scale of Euclid and Roman also demands an increasing reliance on deep learning models for interpreting images (e.g. classification, segmentation, feature extraction for multimodal tasks, etc.) Identifying the concepts learned by our models can inform how we build them and help mitigate the risk of unintended biases or shortcut learning (geirhosShortcutLearningDeep2020, ).
Sparse autoencoders (SAEs) are routinely used for interpreting language models by identifying the set of directions in activation space that can together describe any activation vector in a neural network (makhzani2014ksparseautoencoders, ). Sparsity encourages these directions to be monosemantic, i.e., corresponding to a single concept (elhage2022toymodelssuperposition, ; cunninghamSparseAutoencodersFind2023, ; 2025ApJ…980..183W, ; karvonenSAEBenchComprehensiveBenchmark2025, ). Here, we apply SAEs to reveal the concepts learned by galaxy morphology models.
We construct embeddings for galaxies in the Euclid Q1 data release (collaborationEuclidQuickData2025, ) and then compress those embeddings via SAEs and PCA to identify learned concepts. We show that SAEs efficiently extract candidate morphological features from pretrained networks, comparing against a PCA baseline. For both supervised and self-supervised embeddings, SAEs identify interpretable features aligned with, and beyond the Galaxy Zoo decision trees, while PCA primarily aligns with existing categories.
2 Representations of Galaxy Morphology
Euclid Imaging and Embeddings. We analyze two types of neural network embeddings for galaxy images in the Euclid Q1 dataset. For supervised embeddings, we use the Zoobot models presented in euclidcollaborationEuclidQuickData2025b . These models are pretrained on approximately 1M volunteer-annotated galaxies from four other telescopes and then finetuned using 170k Euclid galaxies with new annotations from Galaxy Zoo Euclid (GZ). Our chosen model, ConvNeXT-Nano, has an encoder with 640-dimensional features.
For self-supervised embeddings, we train a masked autoencoder (MAE, (heMaskedAutoencodersAre2022, )) on 3M Euclid images drawn from the internal Euclid dataset “RR2”. This is $5-10\times$ larger than recent work applying MAE to astronomy images (fathkouhiAstroMAERedshiftPrediction2024, ) and applying other self-supervised learning approaches to Euclid images (euclidcollaborationEuclidQuickData2025, ).
Our MAE is a ViT-S [^1] https://huggingface.co/timm/vit_small_patch16_224.augreg_in21k encoder (30.1M parameters, 384-dimensional features) plus a 3-layer decoder, trained for 2 A100-weeks. We make several design changes compared to the standard ConvNeXT/MAE approach in order to achieve stronger reconstruction performance:
We use a high masking fraction of 90% (vs. 75% as published). We find that networks of just 1M parameters perform well with conventional masking ratios, suggesting that astronomy images have high redundancy. Work using MAE for video (which is intrinsically high redundancy) shows that higher masking ratios are required to force the network to do more than interpolate (feichtenhoferMaskedAutoencodersSpatiotemporal2022, ).
Figure 1: True images (top row), masked inputs (middle row), and reconstructed outputs (bottom row), grouped as follows. Top: challenging images with successful MAE reconstructions. Lower left: images with missing local details (e.g., small starforming clumps, multiple background sources with marginal separation). These cannot be reconstructed if not at least partially included in an unmasked patch. Masking is only applied during training and so these details are still in the embeddings. Lower right: random images (first five) demonstrating that near-perfect reconstructions are the norm.
We use a small patch area of $\times8$ pixels (vs. $\times16$ as published). Astronomy images are relatively sparse, often with large ‘blank’ areas and compact background/foreground galaxies. Random sampling with large patches misses relevant areas entirely, making prediction impossible.
We adjust the frequency of the positional encoding to indicate the angular scale of each image. This is an astronomical analogy to work using MAE for remote sensing (reedScaleMAEScaleAwareMasked2023, ), who adjust the frequency to indicate the ground scale of each satellite image.
Our trained MAE reconstructs masked galaxy features more accurately than professional astronomers. Figure 1 shows our MAE infilling 90% masked galaxy images with near-perfect accuracy. An informal survey of our human colleagues suggests that professional astronomers cannot correctly describe the same masked images. Superhuman reconstruction performance indicates that our MAE extracts meaningful features. We also hope that our MAE will be useful for tasks like patch-level anomaly search (find me images that include patches like these) and targeted infill (remove this small artifact). We release our MAE model on HuggingFace at https://huggingface.co/mwalmsley/models.
Figure 2: Top 10 examples for first five features extracted via PCA (upper), and SAEs (lower), for embeddings extracted from supervised (left) and self-supervised (right) models.
Extracting Features from Embeddings. SAEs learn sparse representations of the embeddings using an overcomplete basis. We train a Matryoshka Sparse Autoencoder (SAE; (bussmann2025learningmultilevelfeaturesmatryoshka, )) with hierarchical group sizes 64, 64, 128, $\cdots$, 1024, batch top-$k$ sparsity ($k=64$), a small L1 penalty, and auxiliary loss for reviving dead neurons (following (2024arXiv240800657O, )). For both self-supervised and supervised embeddings, we train for 200 epochs. The code is publicly available at https://github.com/jwuphysics/euclid-galaxy-morphology-saes.
We use top $k=64$ components in terms of activation frequency. Figure 2 shows the first 5 for each. They explain $\sim 83\%$ of the variance for Zoobot and $\sim 90\%$ of the variance for the MAE). We also perform PCA decomposition of Zoobot and MAE embeddings; the cumulative explained variance using first 64 components is $\sim 90\%$ for Zoobot and $\sim 95\%$ for the MAE.
Interpreting SAE Features via Alignment with Galaxy Zoo. For each learned feature, we compute Spearman rank correlations ($r$) with all Galaxy Zoo morphology labels and take the maximum absolute correlation. [^2] See Appendix B for a similar analysis on whether SAE features are predictable from GZ vote fractions. Higher values indicate stronger correlation with known morphological classifications. We report the mean over the top $k$ PCA or SAE features, where “top” is defined as the highest rank-ordered eigenvalues for PCA, and the most frequently activated neurons for SAEs.
3 Results
Figure 3: Top SAE features (red) are generally more aligned with GZ morphological classifications than PCA (blue) for both the supervised model embeddings (left) and self-supervised model embeddings (right). Both panels show the distribution of maximum Spearman rank correlation $r$ values between top 64 features and any GZ class.
Table 1: Alignment between SAE/PCA Features and Galaxy Zoo vote fractions. We report alignment with “all” GZ classes, as well as the alignment with GZ classes beyond the most common classes (“secondary”), for the top $k=5$ and $k=64$ features extracted via PCA and SAEs. Random feature statistics are described in Appendix C.
Features Supervised Self-supervised
(all) (secondary) (all) (secondary)
PCA ($k=5$) $\bm{0.618\pm 0.149}$ $\bm{0.582\pm 0.097}$ $0.455\pm 0.199$ $0.444\pm 0.193$
SAE ($k=5$) $0.402\pm 0.154$ $0.371\pm 0.161$ $\bm{0.589\pm 0.067}$ $\bm{0.589\pm 0.067}$
PCA ($k=64$) $0.176\pm 0.138$ $0.166\pm 0.137$ $0.434\pm 0.174$ $0.396\pm 0.178$
SAE ($k=64$) $\bm{0.296\pm 0.129}$ $\bm{0.283\pm 0.126}$ $\bm{0.523\pm 0.123}$ $\bm{0.509\pm 0.128}$
Random $041\pm 0.002 0.0$$40\pm 0.002 0.0$$072\pm 0.0003 0.0$$0071\pm 0.0003$
In Figure 3, we show how well the SAE and PCA features correlate with GZ features. An extension of this analysis beyond the top $k=64$ features is presented in the Appendix A (see, e.g., Figure 4).
Supervised: The top $k=5$ principal components of the supervised embeddings strongly correlate with GZ features, even more so than the corresponding top 5 SAE features (see Table 1). When we extend to the top $k=64$, higher-order PCA features show weaker alignment with GZ classes than the SAE features. At the least, this implies that SAE features retain coherency while PCA features become more noisy at $k=64$. We see SAE features outside of the GZ decision tree, e.g., dust lanes in edge-on disk galaxies, elliptical galaxies with bluer companions, etc.
All of the strongest correlations between PCA and SAE features are due to the smooth-or-featured-euclid_* node in the GZ decision tree. This is unsurprising, as the root node classification routes downstream morphological classifications into “secondary” categories. Thus we repeat our alignment tests by omitting this root node, and only using secondary. Overall, we find that alignment with GZ decreases, but we observe the same trends as before.
Self-supervised: The MAE embeddings exhibit surprisingly strong correlations with GZ—they are more correlated than the Zoobot embeddings. We suspect this arises because the MAE embeddings are intent on reconstructing imaging artifacts such as saturated stars or ghosts, which can be diverse and span many pixels. The SAE and PCA features primarily activate on these artifacts, and indeed we see that the most of the strongly aligned GZ classes are problem_* or artifact_*. [^3] Another possibility is that the supervised optimization objective (i.e. aligning to GZ) can (counterintuitively) cram class-separating features into a low-dimensional subspace (e.g., grigg2021selfsupervisedsupervisedmethodslearn, ). In other words, the supervised objective can cause embeddings to span a small number of known, dominant morphological classes, whereas the self-supervised case allows representations to disentangle and occupy a broader space of potentially novel features.
4 Discussion
Limitations. Unfortunately, some SAE features remain difficult to interpret, despite the benefits of using Matryoshka SAEs to combat feature splitting and absorption (Chanin2024AIF, ; bussmann2025learningmultilevelfeaturesmatryoshka, ). Still, SAEs surpass our PCA baseline approach. PCA features become unidentifiable after $\sim 30$ features and show no alignment with GZ. In contrast, our SAE delivers an order of magnitude more aligned features (see Appendix A for more details).
Scientific Discovery. SAEs can extract interpretable features from both supervised and self-supervised models at scale (gao2024scalingevaluatingsparseautoencoders, ). Sparse activations are steerable and can be readily interpreted through automated means (cunninghamSparseAutoencodersFind2023, ; 2024arXiv240800657O, ; ye2024steering, ). Although there are concerns about whether SAEs are effective at learning known concepts, they excel at identifying new features compared to other methods (muhamed2024decodingdarkmatterspecialized, ; movva2025sparseautoencodershypothesisgeneration, ; marks2025auditinglanguagemodelshidden, ; peng2025usesparseautoencodersdiscover, ). Thus, SAEs have tremendous potential to aid scientific discovery by categorizing and sifting through next-generation astronomical datasets (see also, e.g., (2021A&C….3600481L, ; oryan2025identifyingastrophysicalanomalies996, )). By extracting features from self-supervised embedding models, SAEs can act as scalable discovery engines that surface rare or anomalous astrophysical phenomena for future targeted investigation.
Appendix A Interpreting SAE features beyond the top 64
In Figure 4, we show that SAE features beyond the top 64 are still aligned with GZ. For the supervised case (left-hand panel), the correlation grows weaker as we reach the later groups in the Matryoshka SAE: typical correlations for the top 64 features are $\sim 0.3$, whereas mean Spearman correlations in the final group with GZ are near $\sim 0.1$.
Meanwhile, the SAE features for self-supervised embeddings (right-hand panel) all have a moderately high alignment with GZ features. Intriguingly, this does not vary much with the Matryoshka groups, suggesting that the basis set of MAE features is independent of GZ.
We also show PCA again, which exhibits significantly lower alignment with GZ than the SAE features in both the supervised and self-supervised cases first principal component, which has maximum $|r|\approx 0.87$ and separates smooth galaxies from disk or featured galaxies). Principal components beyond the top 64 have nearly zero max Spearman correlation $r$ with GZ features (and are consistent with random), so we omit them from Figure 4.
Figure 4: How aligned are features with GZ, just like in Figure 3? We show results for both supervised (left) and self-supervised (right) approaches. This time we use all of the SAE activations, binned by their Matryoshka group sizes (progressing from lighter to darker red). In all panels, the top 64 elements from PCA are shown in a thin dashed blue line.
Appendix B Feature Predictability
In the same spirit as the “Feature Alignment” analysis, we can also fit linear regression to predict each learned feature from all Galaxy Zoo labels and measure $R^{2}$. This is not a trivial deprojection for the supervised case because the GZ model has a non-linear classification head.
The $R^{2}$ metric probes how predictable any SAE feature is, based on known GZ features. If we are unable to explain the feature variance using existing morphology classifications, then it may suggest that the feature is novel. (It could also signify that the feature is highly polysemantic!)
In Figure 5, we show how well GZ classes can predict the strengths of SAE activations and principal components. We find qualitatively similar results to before: for the supervised case, the PCA features are generally difficult to predict (aside from the first handful), and the SAE activations also become harder to predict as we progress to later Matryoshka groups. For the MAE embeddings, the SAE features are moderately predictable (and PCA features less so). Again, we suspect this is because MAE embeddings contain salient information about image artifacts, and GZ vote fractions can partially reconstruct these features.
Figure 5: How unpredictable—or perhaps, “novel”—are our features via linear combinations of GZ vote fractions? Figure in same style as Figure 4.
Appendix C Expected Spearman Correlations with Random Features
We can fashion a null test by sampling random features and computing the expectation value of the maximum Spearman $r$ by using extreme value statistics: $\mathbb{E}[{\rm max}(r)]\approx\sqrt{2\log(2c)/(n-1)}$, where we have $c=46$ GZ classes, $k=64$ random features, and $n=2378$ or $76061$ test set samples respectively in the supervised and self-supervised test sets. This approximation is valid in the limit of moderate to large $n$, and we use $2c$ rather than $c$ because we take the absolute Spearman correlation (probing a two-tailed distribution). The standard deviation of the maximum Spearman rank can be approximated via $\pi/\sqrt{k(n-1)\log(2c)}$. These expectation values approximate to $0.041\pm 0.002$ for the supervised case, and $0.0072\pm 0.0003$ for self-supervised model embeddings.